OpenStudy (anonymous):

Using distance Formula (-3, 22) and (-14, 35) A. 4 B. SQRT(61) C.SQRT(290) D. 5 E. SQRT(173) F. SQRT(180)

hartnn (hartnn):

need i write the formula again ?

hartnn (hartnn):

what answer are u getting ?

hartnn (hartnn):

Distance between points (x1,y1) and (x2,y2) is \(\huge d=\sqrt{(x_1-x_2)^2+(y_1-y_2)^2}\)

hartnn (hartnn):

x1-x2= .. ? y1-y2=... ?

OpenStudy (anonymous):

(-14-(-3)) sqrt 2+ (35 - 22)) sqrt2

hartnn (hartnn):

then ?

OpenStudy (anonymous):